OpenStudy (anonymous):

Double integration square root 1+2x

OpenStudy (anonymous):

the question is incomplete

OpenStudy (irishboy123):

need to specify double integration wrt what ? and of what?

OpenStudy (anonymous):

Double integration Square root 1+2x dA;R : 0<x<2; 0<y<1

OpenStudy (astrophysics):

\[\int\limits_{0}^{2} \int\limits_{0}^{1} \sqrt{1+2x} dydx\]

OpenStudy (astrophysics):

What's the problem?

OpenStudy (anonymous):

Ya exactly tis is the question,how to solve?

OpenStudy (anonymous):

ok. you have to integrate with respect to y first before x

OpenStudy (anonymous):

I already did it, but I having problem wen integrate dx

OpenStudy (anonymous):